OpenStudy (yuliacosta111):

Write an expression that is equivalent to two over three (4x + 9). eight over threex + 6 eight over threex + 9 two over threex + 6 two over threex + 9

OpenStudy (alexandervonhumboldt2):

2/3*(4x+9)=distribute=8/3x+18/3=8/3x+6